Freedom Cheney Bank started its first day of operations with $10 million in capital. A total of $100

million in checkable deposits is received. The bank makes a $20 million commercial loan and lends

another $30 million in mortgage loans. If required reserves are 5%, what does the bank balance sheet look

like?
Answer #1

Required reserves ($M) = Checkable deposit x Reserve ratio = 100 x 5% = 5

Excess reserves before loans ($M) = Checkable deposit - Required reserves = 100 - 5 = 95

Excess reserves before loans ($M) = Checkable deposit - Required reserves - Total loan = 95 - 20 - 30 = 45

Balance sheet as follows.

ASSETS $M Liabilities and Equity $M
Reserves: Checkable deposits 100
Required reserves 5 Capital 10
Excess reserves 45
Total reserves 50
Loans:
Commercial loan 20
Mortgage loan 30
Total loans 50
Other assets 10
TOTAL 110 TOTAL 110